OpenStudy (anonymous):

How do you find solutions for this linear function A(x)=9x^2-25/3x +5?

hartnn (hartnn):

can u factorize 9x^2-25 ? do u realise its the difference of squares?

OpenStudy (anonymous):

I'm really not sure what I'm doing I don't remember any of this

hartnn (hartnn):

u remember/note this : \(\huge a^2-b^2 = (a+b)(a-b)\) now write \(\huge 9x^2 = (3x)^2 \) and how will be 25 written as ?

OpenStudy (anonymous):

25^2=5^2?

hartnn (hartnn):

yes, correct! now apply that formula and tell me what u get on right side ?

OpenStudy (anonymous):

25^2-5^2=625-25=600?

hartnn (hartnn):

where does 25 come from ?? \(\huge (3x)^2-5^2 = (...)(...)\) how will u fill up those brackets using the formula i gave

OpenStudy (anonymous):

(3x^2+5^2) and (3x^2-5^2)?

hartnn (hartnn):

no! write a= 3x and b=5 now tell me what u get

OpenStudy (anonymous):

Oh you don't put the squares?

hartnn (hartnn):

no, the right side of formula does not have squares, thats y don't put

OpenStudy (anonymous):

Oh okay sorry I'm really bad at this stuff I'm suppose to do the (a+b) (a-b) though right?

hartnn (hartnn):

yes

OpenStudy (anonymous):

So it's just (3+5) and (3-5)?

hartnn (hartnn):

where did x disappear ? it will be \(\huge (3x)^2-5^2= (3x+5)(3x-5)\) got this ?

hartnn (hartnn):

then simplifying this will be easy \(\large A(x)=\frac{ A(x)=9x^2-25}{3x+5}=\frac{(3x+5)(3x-5)}{3x+5}=?\)